Solve using the quadratic equation:
x^2 - 3x - 23 = 5
3x^2 - 2x = 15x - 10

Answer by checkley71(8403)

X~2-3X-23=5 OR X~2-3X-28=0 OR (X-7)(X+4)=0 OR X=7 & X=-4
3X~2-2X=15X-10 OR 3X~2-2X+15X+10 OR 3X~2+13X+10=0 OR (X+10)(X+3)=0 OR X=-10 & X=-3